y''+y'-2y=e^x+4sinx+x^2-x operator method differential equations


1
Expert's answer
2021-01-26T07:32:45-0500



y'=Dy


y"=D2y


(D2+D-2)y=ex+4sin(x)+x2-x


Common solution


D2+D-2=0

"D1=(-1+ \sqrt{1+8})/2=1"

"D2=(-1- \sqrt{1+8})/2=-2"


y(x)=c1ex+c2e-2x


Particular solution for ex

Dekx=kekx

D2ekx=k2ekx

k=1


Dex=ex

D2ex=ex

y(x)=ex+ex-2=2ex-2


Particular solution for sin(x)


Dsin(kx)=kcos(kx)


D2sin(kx)=-k2sin(kx)


k=1


D4sin(x)=4cos(x)


D24sin(x)=-16sin(x)


y(x)=-16sin(x)+4cos(x)-2


Particular solution for x2

Dxk=kxk-1

D2xk=k(k-1)xk-2

k=2


Dx2=2x


D2x2=2(2-1)x2-2=2


y(x)=2+2x-2=2x


Particular solution for x


k=1


Dx=1


D2x=0


y(x)=1-2=-1




Answer: y(x)=c1ex+c2e-2x +2ex-2-16sin(x)+4cos(x)-2+2x-1=c1ex+c2e-2x+2ex-16sin(x)+4cos(x)+2x-5